### v1

# DEPTH: Discourse Education through Pre-Training Hierarchically

Abstract
Language Models (LMs) often struggle with linguistic understanding at the discourse level, even though discourse patterns such as coherence, cohesion, and narrative flow are prevalent in their pre-training data. Current methods address these challenges only after the pre-training phase, relying on expensive human annotated data to align the model. To improve the discourse capabilities of LMs already at the pre-training stage, we introduce DEPTH, an encoder-decoder model that learns to represent sentences using a discourse-oriented pre-training objective. DEPTH combines hierarchical sentence representations with two objectives: (1) Sentence Un-Shuffling, and (2) Span-Corruption. This approach trains the model to represent both sub-word-level and sentence-level dependencies over a massive amount of unstructured text. When trained either from scratch or continuing from a pre-trained T5 checkpoint, DEPTH learns semantic and discourse-level representations faster than T5, outperforming it in span-corruption loss despite the additional sentence-un-shuffling objective. Evaluations on the GLUE, DiscoEval, and NI benchmarks demonstrate DEPTH’s ability to quickly learn diverse downstream tasks, which require syntactic, semantic, and discourse capabilities. Overall, our approach extends the discourse capabilities of T5, while minimally impacting other natural language understanding (NLU) capabilities in the resulting LM. [^1] We release our repository for pre-training and fine-tuning encoder-decoder models as well as the checkpoints we used for this study.
1 Introduction
Discourse understanding—the ability to understand how sentences and broader textual units form cohesive narratives—is fundamental to effective communication. However, LMs often struggle with this, especially when dealing with long and complex inputs. This limitation hinders the performance of LMs on various natural language processing (NLP) tasks such as persuasive argumentation [16], summarization [69], essay scoring [36], dialogue systems [20], and following intricate instructions [60]. Existing LMs frequently generate text that lacks compositionality, suffers from disconnected ideas and contradictions, or fails to develop core ideas across sentences [17, 34, 35]. These issues stem from their limited ability to capture the underlying structure, implied meanings, and relationships between ideas that are crucial for effective communication.
Early attempts to incorporate discourse awareness into pre-training, such as Next Sentence Prediction (NSP) in BERT [12] and Sentence Order Prediction (SOP) in ALBERT [26], proved overly simplistic and hindered learning effective discourse representations [32, 26, 45]. Subsequent models like Sentence-level Language Model (SLM) [27] and CONPONO [21] improved discourse capabilities in LMs but lacked generative capabilities. On the other hand, the next-token prediction pre-training objective provided LMs like Generative Pre-trained Transformer (GPT) [43, 44, 4, 40] with powerful generative capabilities. However, without a dedicated and costly alignment phase [42], these LMs tend to falter at discourse-level evaluation benchmarks, even when scaled to over 100 billion parameters [35, 34, 5]. We therefore focus our studies on encoder-decoder models such as Text-to-Text Transfer Transformer (T5) [45], which can compete with (and sometimes outperform) GPT-3 [4] on long (e.g., paragraph-level) discourse-oriented tasks even at $\approx\frac{1}{400}$ the size [35, 34]. [^2] The T5 model used in this study, released by Wei et al. [60], was fine-tuned to follow instructions.
To improve the discourse-capabilities of encoder-decoder models, we propose DEPTH (Discourse-Education through Pre-Training Hierarchically), a hierarchical language model that learns representations for both sub-word and sentence-level tokens. DEPTH builds on the architecture of T5 and incorporates the discourse-oriented objective of SLM. By employing a hierarchical attention mechanism across sub-word and sentence level tokens, DEPTH captures both fine-grained semantic dependencies and broader inter-sentence relationships. Furthermore, its simplified loss function unifies a discourse-oriented pre-training objective, previously limited to encoders, with the traditional teacher-forcing optimization used in decoders. When randomly initialized, our DEPTH model obtains meaningful representations for downstream tasks much faster than our baseline T5. Continuously pre-training T5 models with the DEPTH objective improves the discourse capabilities of the resulting models, without sacrificing performance in downstream NLU tasks.
2 Method
Pre-training DEPTH involves simultaneously performing span corruption [45], while also un-shuffling sentences as in Lee et al. [27]. In Section 2.1 we introduce a new tokenizer for DEPTH, which combines T5’s tokenizer with the sentence segmentation operation required for sentence un-shuffling. In Section 2.2, we detail how we combined the pre-training objectives of both models into the sequence-to-sequence framework of T5. Next, in Section 2.3, we discuss how we introduce hierarchical representations during pre-training, and how this hierarchy encourages the model to learn discourse representations. Finally, in Section 2.4, we demonstrate how to combine the losses of T5 and SLM into a unified objective that is conducive to traditional teacher-forcing.
2.1 Tokenization
We introduce a tokenization function, $t$, which transforms an input string, $s$, into a sequence of tokens, $X=(x_{1,1},x_{1,2},\ldots,x_{m,len(m)})$, in our vocabulary, $V$. In this formulation, each token $x_{i,j}$ denotes the $j$’th token of the $i$’th sentence, $m$ represents the number of sentences, and $len(i)$ represents the length of the $i$’th sentence. $V$ also includes special tokens <EOS>, <BOS>, <PAD>, and a set of sentinel tokens, $V_{sentinel}$, of the form <special_token_z> as in the original T5 paper.
Figure 1: Given an input document, like “Hello! How are you?”, DEPTH introduces sentence tokens (<SENT_i> and <EOSEN>) as part of tokenization. Then, as part of corruption, DEPTH applies both span masking (in this case, applying a single masked span over the tokens “are” and “you?”), and shuffles the sentences in the encoder with probability $0.5$. As part of training, sentence tokens in the encoder can only attend to tokens in their own sentence. Sentence tokens in the decoder can only attend to earlier sentence tokens (both in the encoder and in the decoder). Attention among sentence tokens is represented as arrows, where the source is the query token, and the target is the key token (cross attention uses dotted arrows, while self attention uses full arrows).
To facilitate DEPTH’s hierarchical structure, we segment sentences using NLTK [3] (as in SLM) in each example, and create $k+1$ new tokens (where $k$ is the maximum number of sentence tokens per example [^3] We follow Lee et al. [27] by using $k=20$ sentence tokens.):
[…]
We augment our tokenizer function $t$ to form $t^{\prime}$, which maps sequence $s$ to tokens in $V^{\prime}$. In each sentence, we prepend a <SENT_i> token and append a <EOSEN> token as follows:
[…]
The integer $i$ within <SENT_i> represents a sentence’s index, sampled uniformly at random from ${1,...,k}$. We sample sentence token IDs without replacement to eliminate information about the sentence’s position. We truncate sentences beyond the $k$’th to limit our vocabulary size.
Unlike the original SLM model (an encoder-only LM with an auxiliary pointer-decoder for pre-training), DEPTH is an encoder-decoder and cannot directly “point” to the latent vector of the next sentence. We address this by explicitly introducing <SENT_i> and <EOSEN> tokens into the vocabulary for the model to predict their IDs. Notably, we introduced <EOSEN> to signal that the next token is some <SENT_i> or the end of the example (<EOS>). In essence, <EOSEN> acts as a “signal” token so that the model can dynamically adjust its attention mask in the decoder (i.e., “should it attend to all previous tokens, or just to previous sentence tokens?”).
Formally, we define a pre-tokenization function $f$, which maps $s$ to $s^{\prime}$, where $s^{\prime}$ includes <SENT_i> and <EOSEN> tokens: $f:s\rightarrow s^{\prime}$. The tokenized input for DEPTH is produced with the function $T$, such that $T(s)=t(f(s))=X$. In practice, we use the SentencePiece [25] tokenizer as our function $t$, and adjust it to support DEPTH’s additional pre-training objective.
2.2 Corruption
Span-Masking:
We apply a corruption process to each batch of tokenized sequences. We sample masked spans using a geometric distribution (as in Joshi et al. [23] and Raffel et al. [45]), parameterized with an average span length of $\lambda$ and a masking probability of $p$. Spans overlapping with sentence tokens (<EOSEN> or <SENT_i>) are ignored. Masked token spans are replaced with a single sentinel token <special_token_z>, where ($z$) is a uniformly sampled integer from ${0,\ldots,99}$. The missing token sequences appear after the corresponding sentinel token in the labels. Note that in the original T5 implementation, sentinel tokens appear in the same incrementally decreasing order in each example (<special_token_99> followed by <special_token_98>, etc…). We randomly sample sentinel token ID’s for DEPTH to eliminate hints about sentence positions from the sentinel tokens. For example, with the T5 scheme for span-masking, the presence of <special_token_99> indicates that the sentence in which it appears comes first, making sentence un-shuffling too easy.
Sentence Un-Shuffling:
Given an input sequence of up to $k$ sentences, we randomly shuffle the order of sentences in the model input (but do not shuffle the order of tokens within a sentence). We then task the model with reconstructing the original order of the sentence tokens in the target. We shuffle all examples in a batch with probability $p$ (we use $p=0.5$, as in SLM). We use this objective to encourage DEPTH to attend to other sentences for discourse-level cues, rather than relying on local words and phrases. By disrupting the original sentence order, DEPTH encourages learning of (1) the complete meaning of individual sentences, independent of their surrounding context, and (2) representations that encode how sentences relate to each other (e.g., discourse markers, co-reference, entailment). We show the full DEPTH corruption process in figure 1.
2.3 Attention Masks
Our baseline model (T5) utilizes bidirectional attention in the encoder, auto-regressive attention in the decoder, and full cross attention from the decoder to the encoder. However, T5’s formulation does not account for the hierarchical treatment of sentences used by SLM and DEPTH.
We define non-sentence tokens, $x_{reg}$, as tokens $x\in X$ where $x\notin S$. We compose attention masks to impose hierarchy. As part of encoder self-attention, non-sentence tokens can attend to all other tokens in the corrupted input sequence, while <SENT_i> tokens can only attend to tokens within their own sentence (including sentinel tokens). As part of decoder self-attention, all tokens have an auto-regressive attention mask, but <SENT_i> tokens can only attend to past sentence tokens. Finally, as part of cross attention, non-sentence tokens in the decoder can attend to the entire input in the encoder, while sentence tokens in the decoder can only attend to sentence tokens in the encoder. This scheme is depicted in figure 1.
This scheme encourages the model to rely on sentence token representations in the encoder to predict the correct next sentence token in the target. Our approach encourages sentence tokens in the input to encode discourse information (about the relative position of the sentence) that is useful for predicting the next sentence token via cross attention. As a second-order effect, this encourages non-sentence tokens in the encoder to provide their respective sentence token with useful discourse-level information.
2.4 Loss Formulation
Given that $\hat{S}$ is the set of all sentence tokens within our target sequence, let $Y=\{y_{1,1},y_{1,2},\ldots y_{m,len(m)}\}$ be the unified target sequence with tokens that can be either part of the span-masking task (non-sentence tokens) or the un-shuffling task (sentence tokens). The prediction $\hat{Y}=\{\hat{y}_{1,1},\hat{y}_{1,2},\ldots\hat{y}_{m,len(m)}\}$ represents the model’s output for each token in $Y$.
The loss for the DEPTH model, considering both reconstruction (a.k.a., span-masking) and sentence un-shuffling tasks, is simply the cross-entropy loss, which is calculated as:
[…]
In this formulation, the summation runs over all sentences in the input (up to $k$), and each token within each sentence (up to $len(i)$, where $i$ is the index of some sentence) in the target sequence $Y$, thus allowing us to decompose DEPTH’s effectiveness in learning each pre-training task. We explore additional loss formulations in Appendix B.1.
3 Experimental Setup
The aim of our experiments is to measure the effectiveness of DEPTH against a standard encoder-decoder model. Accordingly, our experiments explore the learning dynamics of DEPTH model relative to a T5-Base (220M parameter) baseline. We pre-train both models on the C4 dataset [45, 14] to resemble the manner in which the original T5 was trained (see additional reasoning in Appendix D.2). We chose to use Base-sized models given limited computational resources, and ease of reproducibility (following the example of SLM, PMI, PEGASUS, and T5 [27, 28, 67, 45]). We did not use SLM as our baseline since its codebase and checkpoints are not released, and it cannot perform free-form text generation.
We chose to run our experiments without example packing since this is how the SLM model, which inspired DEPTH, was trained. Furthermore, example-packing in T5 enables unrelated examples to impact the model’s decisions, thereby harming performance [24, 49]. While example-packing is critical for more computationally-efficient training [13], we were interested in measuring which model was able to use training examples more effectively. We examine the impacts of avoiding example packing in Appendix A.2.
[…]
We conduct our pre-training experiments under two categories:
  1. 1.
    From Scratch (FS): Both T5 and DEPTH models are randomly initialized, and pre-trained on C4 with their respective objectives.
  2. 2.
    Continuous Pre-Training (CPT): Both T5 and DEPTH models are initialized from the T5-Base checkpoint on HuggingFace [63], and continue to pre-train on C4 with their respective objectives.
We note that our CPT experiments build on top of T5 models that have been trained for over 1T tokens, whereas the amount of tokens they see during continuous pre-training is relatively minuscule (67$\approx$ 67$\times$ 67 × fewer tokens for T5, and 80$\approx$ 80$\times$ 80 × fewer tokens for DEPTH). For reference, we show the relative pre-training configurations of similar models in Appendix D.1.
3.1 Fine-Tuning Experiments
We follow up our pre-training experiments with Natural Language Understanding (NLU) downstream tasks. We evaluate our models on natural language inference (MNLI, [62]), sentiment analysis (SST2, [50]), and grammar (CoLA, [59]) within the GLUE benchmark [58]. We also use the DiscoEval suite [5] to evaluate models on their understanding of discourse. We use two tasks from this benchmark: Sentence Permutation (SP) and Discourse Coherence (DC). SP involves shuffling sentences, and predicting which of the shuffled sentences originally appeared first, while DC involves predicting whether or not the input paragraph was coherent. Finally, we measure our model’s generative abilities on the Natural Instructions (NI) dataset [37], which measures the ability of LMs to follow instructions, and served as a benchmark for NanoT5 [39].
Our experimental framework is inspired by Pythia [2], as we evaluate the performance of our models on downstream tasks from intermediate checkpoints. Unlike Pythia, we run evaluation with checkpoints from both T5 and DEPTH models, gathered at steps $\{2K,4K,\ldots,512K,1M\}$ in order to examine these models’ emergent capabilities. The exponential distance between these checkpoints allows us to scale intermediate checkpoint evaluation to much longer training runs  [^5] Evaluating intermediate checkpoint performance every 10,000 steps (as was done in Pythia) on datasets as large as MNLI is unfeasible with our limited computational resources..
4 Results
4.1 C4 Pre-Training
During pre-training, we find that DEPTH consistently achieves a lower validation loss than a comparably trained T5 model. This is true for both FS pre-training, and CPT pre-training. Furthermore, when we isolate the reconstruction loss (the same objective used by T5, not including sentence tokens), we find that DEPTH outperforms T5 despite balancing an additional pre-training objective (Figure 3 for FS and Figure 3 for CPT). These results are consistent with the findings in SLM, where their model converged faster, and on fewer tokens than models such as BERT and T5.
Figure 2: From Scratch Pre-Training loss (validation) for both T5 and DEPTH
Figure 3: Continuous Pre-Training loss (validation) for both T5 and DEPTH.
While we were not able to match the performance of the baseline model of Raffel et al. [45] (see Appendix A for speculations on why), we have obtained the lowest loss scores among PyTorch implementations of T5 models. Specifically, in our FS setting, we find that our randomly initialized T5 model outperforms the validation loss of NanoT5, achieving $1.65$ vs. $1.95$ at step 64,000.
4.2 GLUE Fine-Tuning
We found that over the course of FS pre-training, both models improved on GLUE tasks. However, T5’s improvement pattern was sporadic and slow when compared to that of DEPTH (top row of Figure 4). We found it difficult to replicate the results of the original T5 (both on GLUE tasks and the pre-training loss) as discussed in Appendix A. We project that with more substantial training (i.e., 1-3 million pre-training steps, and a larger batch size of $\geq 2048$ examples per batch, as in T5 [45]), DEPTH could match or exceed the performance of T5 and SLM on downstream tasks.
(a) CoLA FS
(b) MNLI FS
(c) SST-2 FS
(d) CoLA CPT
(e) MNLI CPT
(f) SST-2 CPT
Figure 4: GLUE results for models trained from scratch (FS) and from pretrained (CPT). Top row: From Scratch (FS), Bottom row: From Pretrained (CPT).
In the CPT setting (Figure 4, bottom row), we found that DEPTH and T5 perform similarly, both improving only slightly beyond the baseline. Initially, DEPTH seemed to perform worse than the baseline model, and did consistently worse than the T5-CPT model until later stages of pre-training. We speculate that this dip in performance is related to the change from the learned objective of the T5 model (from which DEPTH-CPT was initialized) from span-masking to span-masking and sentence un-shuffling. We share our full results on GLUE in Appendix E.
4.3 DiscoEval Fine-Tuning
We find that in the FS setting DEPTH consistently outperforms T5 across DC tasks, indicating its robustness in understanding discourse (Figure 5, top row). This suggests DEPTH’s pre-training objective is particularly beneficial for tasks that require a deep understanding of narrative structures (both in conversations as in DC-Chat, and more formal and informative texts as in DC-Wiki). We note that between steps 32k and 64k, DEPTH experienced a large positive boost in performance on DC-Wiki, perhaps indicative of an emergence [61] of discourse understanding during this phase of pre-training. For T5, we found that the model struggled to learn the SP-Arxiv task, achieving random-guess accuracy late in pre-training. However, in SP-Wiki and SP-Rocstory, T5 improves in performances between steps 64k and 128k, perhaps indicating an emergent ability occurring within this timeframe. We report our full results on DiscoEval in Appendix F.
(a) Sentence Permutation (SP) FS
(b) Discourse Coherence (DC) FS
(c) Sentence Permutation (SP) CPT
(d) Discourse Coherence (DC) CPT
Figure 5: DiscoEval benchmark results for DEPTH and T5 models. Top row: From Scratch (FS), Bottom row: From Pretrained (CPT).
Table 1: Comparison of various models on the SP and DC tasks within DiscoEval. All models aside from T5 and DEPTH and encoder-only models trained with discourse-oriented objectives.
Model SP DC
RoBERTa-Base 38.7 58.4
BERT-Base 53.1 58.9
BERT-Large 53.8 59.6
CONPONO 60.7 72.9
SLM (1M) 72.4 75.4
SLM (3M) 73.4 76.1
T5-Base 58.1 80.5
T5-FS 40.91 63.31
T5-CPT 59.48 82.27
DEPTH-FS 55.45 76.22
DEPTH-CPT 65.59 82.49
While DEPTH outperformed other models in DC tasks, it failed to reach a high performance level on SP tasks (under-performing relative to SLM, as seen in Table 1). This problem stems already from the pre-training stage, where DEPTH’s sentence un-shuffling accuracy is relatively low ($\leq 5\%$ accuracy on shuffled sentence tokens; see Appendix B for additional details). This highlights the complexity of sentence un-shuffling relative to older discourse objectives like NSP and SOP. Surprisingly, while this task was challenging for DEPTH, SLM reported strong performance on sentence un-shuffling. SLM used a dedicated pointer-generator network that consists of a shallow DNN. This module “points” to one of at most $k$ sentences as it iterates over a target sequence consisting of only sentence tokens. Also, SLM’s non-sentence tokens cannot observe sentence-level tokens as part of the reconstruction loss, avoiding a potential “distraction” in their task.
4.4 NI Fine-Tuning
In the FS setting (Figure 6(a)), we observe that DEPTH outperforms T5 in the NI benchmark, with a notable leap in performance between steps 16k and 32k. T5, by comparison, only improves significantly after step 64k, and obtains worse performance than DEPTH by the end of training. However, in the CPT setting (Figure 6(b)), DEPTH’s pre-training appears to hinder downstream performance compared to T5, possibly due to the domain shift from T5’s pre-training task to DEPTH’s pre-training task, which involves learning from shuffled inputs. We present a more complete analysis of these results in Appendix G.
(a) Natural Instructions (NI) FS
(b) Natural Instructions (NI) CPT
Figure 6: NI results for DEPTH and T5 models.
5 Related Work
The potential of encoder-decoder architectures in today’s NLP landscape cannot be overstated. These architectures dominate context-heavy tasks ranging from translation [71, 65, 55] to summarization [67, 19, 55], and even following instructions across diverse domains [1, 60, 8]. Like their decoder-only counterparts, encoder-decoders are able to accommodate long inputs [19], and scale effectively effectively as a function of model size and training data [54]. Recently, Reka AI released a series of encoder-decoder models [41], where their dense 21B parameter model outperformed all models of its size in lmsys benchmark [70]  [^6] This Reka model is competitve with mixtral 8x22b [22] (which was trained with significantly more parameters using a mixture-of-experts architecture)..Encoder-deocder models are also strong multi-modal learners [41, 64, 15, 66]. When scaled sufficiently, encoder-decoders like Reca-Core may be competitive with state of the art models like GPT-4 [40], Gemini [56], and Claude-3.
While specialized encoder-decoder models such as PEGASUS [67], DialogVED [6], and a multi-party dialogue pre-training model [30] demonstrate the value of discourse-oriented tasks for encoder-decoder models, they have limited utility for broader tasks. Long-T5 [19] and UL2 [55] improved the ability of encoder-decoders to handle long contexts, but did not explicitly tackle discourse understanding. Flan-T5 [60] and Ex-T5 demonstrated the applicability of encoder-decoders across a variety of tasks, including ones that are heavily discourse dependent. However, these models depend on a vast yet costly annotated dataset to learn human preferences. The ERNIE series [68, 52, 53] integrates knowledge graphs and heuristic-guided masking for enriched representations but relies on extensive resources and faces challenges in timely knowledge updates. DEPTH, on the other hand, learns discourse through a generalized pre-training objective, and is not task-oriented or annotation-dependent.
6 Limitations
Given our lack of computational resources (Appendix C), we were not able to pre-train our models with a batch size that would allow an aggressive learning rate like that used in Raffel et al. [45]’s T5 (see Appendix A for additional details). We also pre-trained on substantially fewer tokens than T5. As a result, our model converges to a worse loss during pre-training, and performs worse on various downstream tasks. We also lacked sufficient computational resources to compute confidence intervals or statistical significance for our downstream experiments. [^7] Running a single downstream experiment on MNLI takes $5$-$7.5$ hours. We run $\approx 120$ experiments for each of 10 benchmarks, and do not have the capacity to repeat experiments 5$\geq$ 5$\times$ 5 × to obtain statistical significance.
Encoder-decoder LMs have fewer tools available for computationally efficient pre-training. For example, FlashAttention [11, 10], which provides a massive training speedup, is not available for encoder-decoder models. Furthermore, evolving approaches at the software level, such as Multi-Query Attention [47] for faster computation, and Ring-Attention [31] for extending context length, are not thoroughly vetted for encoder-decoder LMs. It is therefore difficult to create scalable pre-training experiments with new encoder-decoder architectures and objectives.
7 Conclusions and Future Work
DEPTH’s new pre-training objective and hierarchical representations complement efforts to scale model size, parallelize architectures, and acquire high quality data for pre-training. Despite training over fewer tokens, DEPTH significantly outperformed T5 both during pre-training and during fine-tuning. DEPTH’s remarkably efficient learning and downstream performance on discourse oriented tasks underscore the importance of discourse-oriented pre-training.
Looking forward, the application of DEPTH to RAG [29], especially over sentence “chunks”, presents an exciting avenue for future research. Additionally, extending DEPTH’s pre-training objectives to encompass higher-level discourse units—such as paragraphs, chapters, and whole documents—offers further flexibility emerging hierarchical RAG systems [7]. Moreover, conducting further experiments with larger DEPTH models is helpful for understanding the scalability of discourse-focused training objectives. Such investigations could reveal whether the promising capabilities observed in DEPTH are amplified with increased model capacity.
Appendix A Challenges in Replicating T5
A.1 Our Speculations
We recognize that both DEPTH and T5 CPT models quickly reach a significantly lower reconstruction loss than their FS counterparts. While the original T5’s reached a final training loss of $\approx 0.75$ (with 512k steps of batch size 128 with packing) [45], our training loss reached $\approx 0.8$ (after 800k steps of batch size 200 without packing). We propose the following ideas to explain this gap:
  1. 1.
    The baseline model from which CPT models are initialized is trained on 1T tokens. This is 20$\approx$ 20$\times$ 20 × greater than the amount of tokens we used to train FS models.
  2. 2.
    Our examples never consist of the later text in long documents. By truncating text after 512 tokens, our FS models might miss out on valuable text to train on.
  3. 3.
    Our training examples consists of 2$\approx$ 2$\times$ 2 × fewer tokens (on average) than examples the original T5 was trained on. Furthermore, there is much greater variance in example lengths in our pre-training experiments. These factors may impact our models’ learning dynamics (e.g., learning effective positional representations given the presence of irregular padding patterns).
  4. 4.
    The T5 baseline plot was reported using a masking probability of $0.15$, which is 2×2$\times$2 × lower than the one we used. Our higher masking probability makes the reconstruction task more challenging.
  5. 5.
    T5 models using the T5x framework use aggressively high learning rates, which can lead to a different, and perhaps more effective training dynamics than the ones we found in our FS experiments. Using such high learning rates in our settings caused our models to diverge.
A.2 Example Packing
Given our choice of avoiding example-packing, we found that we were not able to pre-train T5 with the same hyper-parameters used in [45]. Specifically, we found that with a batch size of 128 and a learning rate of $1e^{-2}$, our model consistently diverged. This issue persisted with a learning rate of $1e^{-3}$. To stabilize our loss given the absence of packing, we used a lower maximum learning rate ($1e^{-4}$), which is in line with those used to pre-train BERT, SLM, and PMI. On the other hand, we see that given the same training parameters (i.e., learning rate and batch size), pre-training with packing can converge at a high learning rate (see Figure 7).
We speculate that packing acts in a similar way to increasing the batch size during training. The model is exposed to loss on a greater amount of tokens in each optimization step, and is therefore able to generalize even with a larger learning rate.
One possible side effect of avoiding example-packing is the truncation of long examples (examples are not dynamically chunked, so every token past the context limit of 512 is ignored). We empirically find that while T5 suffers greatly from no packing, DEPTH is able to train effectively despite these limitations.
Figure 7: Exploration of packing and learning rates when pre-training T5 models. “High LR” corresponds to a learning rate of $1e^{-2}$, while “Low LR” corresponds to a learning rate of $1e^{-4}$.
Appendix B DEPTH Loss Decomposition
When we decompose DEPTH’s losses in the FS setting, we find that sentence loss is consistently lower than reconstruction loss, but plateus early on into training. The overall loss is dominated by the reconstruction loss, reflected by overlapping lines in Figure 9. In the CPT setting (Figure 9), we find that both of DEPTH’s losses plateu sooner, and that the sentence loss is approximately equal (though with higher fluctuations) to the reconstruction loss.
Figure 8: Decomposition of from-scratch pre-training losses (validation) for DEPTH.
Figure 9: Decomposition of continuous pre-training losses (validation) for DEPTH
We note that DEPTH’s loss over sentence tokens in the FC setting, is close to that which the DEPTH-CPT achieved (both in the range of 0.3-0.4, where the more examples are shuffled, the higher the sentence loss). In practice, given comperable ratios of shuffling sentences, CPT DEPTH outperforms FS DEPTH in predicting the next sentence accurately during pre-training ($\approx 1\%-3\%$ higher given a fixed shuffling ratio). We speculate that the better representations for non-sentence tokens in the CPT setting is the reason for this performance boost.
B.1 Weight of Sentence Loss
We explored the impact of increasing the weight of the sentence loss during DEPTH pre-training. Our “Baseline” run is the DEPTH model we reported on in the main body of the paper. In our “Sentence Weight 1x” run, our loss is composed of the average loss over sentence tokens plus the average loss over non-sentence tokens (as opposed to the average loss over all tokens). This formulation places increased weight on sentence tokens, since there are significantly fewer sentence tokens than non-sentence tokens. In the “Sentence Weight 5x” run, we weighed the sentence loss 5×5$\times$5 × more than reconstruction loss.
We found that increasing the weight of this loss had minimal impact on the model’s accuracy in predicting sentence tokens, and adversely harmed the model’s loss (see Figure 10).
Figure 10: We explore the impact of weighing sentence loss more than reconstruction loss, and find that it has minimal impact beyond early stages of pre-training.
Appendix C Computational Resources
We utilize 4 A40 GPUs, and 64 CPUs for training. We use a batch size of 200, since it helps us achieve much better GPU memory utilization. We leverage 16 CPUs for each GPU in order to increase the data loading time to accommodate for DEPTH’s more-complex corruption method, and to allow effective optimizer offloading with DeepSpeed Zero2 [46].
Appendix D Pre-Training Considerations
D.1 Pre-Training Scale
In table 2 below we show the relative magnitude of DEPTH’s pre-training. Specifically, we compare the number of observed tokens and optimization steps of our T5 and DEPTH models to comparably sized models (such as SLM, BERT, and RoBERTa). We highlight that the models we independently pre-trained (bottom 4 rows of the table) observed far fewer tokens than comparable LMs.
Table 2: Hyper-parameters of comparable models to DEPTH. We show published hyper-parameters in the top rows of the table, and the models we train ourselves in the bottom of the table. We mark all models initialized from publically released pre-training models with (*). Note that CONPONO did not report the number of tokens it pre-trained on, so we exclude that value from the table above.
Model Tokens Steps Batch Size # Params Learning Rate
SLM-1M $125B$ $1M$ $256$ $\approx 110M$ $1.5e^{-4}$
SLM-3M $375B$ $3M$ $256$ $\approx 110M$ $1.5e^{-4}$
BERT-Base $137B$ $1M$ $256$ $\approx 110M$ $1e^{-4}$
BERT-Large $137B$ $1M$ $256$ $\approx 340M$ $1e^{-4}$
RoBERTa-Base $2.2T$ $500k$ $8000$ $\approx 110M$ $1e^{-4}$
CONPONO (*) - $256k$ $256$ $\approx 110M$ $1e^{-4}$
T5-Base $1T$ $1M$ $2048$ $\approx 220M$ $1e^{-2}$
T5-FS $58.6B$ $1M$ $200$ $\approx 220M$ $1e^{-4}$
T5-CPT (*) $15B$ $256k$ $200$ $\approx 220M$ $1e^{-4}$
DEPTH-FS $48.9B$ $1M$ $200$ $\approx 220M$ $1e^{-4}$
DEPTH-CPT (*) $12.5B$ $256k$ $200$ $\approx 220M$ $1e^{-4}$
D.2 Pre-Training Data
The rationale behind selecting C4 extends beyond its sheer volume and diversity. Given DEPTH’s architectural roots in the T5 model, which was originally pre-trained on C4, leveraging the same dataset facilitates a direct comparison of the enhancements our model introduces. This baseline compatibility is crucial for isolating the effects of our architectural and methodological innovations on the model’s performance. Furthermore, C4 is a subset of both Dolma [51] and RedPajama [9]. These datasets were used to train the most capable fully-open-source LMs (with released data) to date: OLMO [18] and LLama [57] respectively. This suggests that C4 is an effective component in pre-training effective LMs, and offers multiple additional datasets components that would be new to our model for future training. Finally, Muennighoff et al. [38] suggest that pre-training over the same data up to 4×4$\times$4 × still improves model performance, while [45]’s model has not completed even one full iteration over the C4 dataset. This hints at the viability of continuing to pre-train T5 on the same dataset it was already pre-trained on.
Appendix E GLUE Tables
In this section we show the full results from our downstream experiments on GLUE tasks. Table 3 is reflected in the top row (FS) of Figure 4, while Table 4 is reflected in the bottom row (CPT) of figure 4. Consistent with our hypothesis, we found that both DEPTH and T5 improve across downstream tasks as a function of the pre-training steps they’ve taken. While DEPTH outperforms T5 across all tasks is the FS setting, it did not reach the scores of Raffel et al. [45]’s T5 model. In the CPT setting, T5 and DEPTH perform quite comparably. In fact, in the penultimate checkpoint (128k) we found that DEPTH outperformed T5 on all tasks except for CoLA.
Table 3: GLUE benchmark results for From Scratch (FS). Note that the first two rows are reported by Raffel et al., 2019, while all later rows are the best reported results on the validation set across 3 attempted learning rates.
Model CoLA SST-2 MNLI
Matched Mismatched
T5-Base @ 0 (Test) 12.3 80.62 68.02 68.0
T5-Base @ 1M (Test) 53.84 92.68 84.24 84.57
T5-Base @ 1M (Val) 53.98 94.73 87.28 87.1
T5 @ 2k 8.62 80.08 52.54 52.3
DEPTH @ 2k 8.67 78.52 57.07 58.14
T5 @ 4k 6.23 80.66 53.16 53.89
DEPTH @ 4k 7.75 80.08 59.31 59.62
T5 @ 8k 4.66 82.23 54.33 54.62
DEPTH @ 8k 6.93 79.69 58.92 59.73
T5 @ 16k 8.99 80.96 54.1 54.1
DEPTH @ 16k 10.94 81.25 62.79 61.46
T5 @ 32k 10.72 81.64 55.18 55.6
DEPTH @ 32k 7.73 82.81 71.23 72.7
T5 @ 64k 6.86 82.42 57.68 60.88
DEPTH @ 64k 27.78 86.72 73.84 76.05
T5 @ 128k 12.85 83.2 69.61 69.82
DEPTH @ 128k 38.01 88.87 77.5 78.06
T5 @ 256k 11.78 85.94 72.82 73.39
DEPTH @ 256k 45.57 91.31 79.45 80.07
T5 @ 512k 19.96 86.52 74.22 74.26
DEPTH @ 512k 47.14 91.11 80.42 81.43
T5 @ 1M 29.35 88.77 74.53 75.37
DEPTH @ 1M 45.91 91.41 81.0 81.96
Table 4: GLUE benchmark results for Continuous Pre-Training (CPT). As in the FS setting, we report our results on the validation set after a hyper-parameter sweep over 3 learning rates.
Model CoLA SST-2 MNLI
Matched Mismatched
T5-Base @ 1M (Test) 53.84 92.68 84.24 84.57
T5-Base @ 1M (Val) 53.98 94.73 87.28 87.1
T5 @ 2k 57.41 95.02 86.93 87.06
DEPTH @ 2k 53.93 94.34 86.38 86.2
T5 @ 4k 55.18 95.02 87.4 87.34
DEPTH @ 4k 47.11 94.34 87.14 87.01
T5 @ 8k 55.35 95.21 87.47 87.31
DEPTH @ 8k 50.67 94.43 86.62 86.52
T5 @ 16k 54.75 95.7 86.91 86.67
DEPTH @ 16k 52.65 94.34 86.62 86.67
T5 @ 32k 54.95 95.21 86.64 86.06
DEPTH @ 32k 53.79 94.63 86.95 87.02
T5 @ 64k 54.77 95.21 86.96 86.93
DEPTH @ 64k 52.95 94.14 86.65 86.91
T5 @ 128k 58.62 95.21 86.79 86.67
DEPTH @ 128k 56.21 95.61 87.42 87.64
T5 @ 256k 57.62 95.21 87.27 87.22
DEPTH @ 256k 56.78 95.02 86.86 86.45
Appendix F DiscoEval Tables
In this section we show the full results from our downstream experiments on discourse tasks from the DiscoEval benchmark. In Table 5, we show the full results of our models in the FS setting, while in Table 6 we show the full results of our models in the CPT setting. These tables reflect the top and bottom row of Figure 5 respectively.
In the FS setting, we found that DEPTH’s wins over T5 are even more pronounced in DiscoEval than they were in GLUE. Specifically, T5 struggles to learn discourse tasks (especially SP) during early stages of pre-training. On the other hand, DEPTH was highly effective in discourse tasks already from early pre-training checkpoints. In the CPT setting, we found that DEPTH still outperformed T5, despite the fact that the original checkpoint was pre-trained substantially with a different objective.
Table 5: DiscoEval Downstream Full Training (FS) Results
Model SP DC
Arxiv Wiki Rocstory Chat Wiki
Baseline T5 @ 1M 52.76 51.07 70.58 68.99 92.09
T5 @ 2k 21.0 20.9 21.2 55.18 53.59
DEPTH @ 2k 34.81 40.19 51.51 55.27 53.2
T5 @ 4k 20.4 21.4 20.9 57.18 55.37
DEPTH @ 4k 35.72 40.38 52.03 56.49 54.74
T5 @ 8k 20.6 20.9 21.0 57.42 54.57
DEPTH @ 8k 36.43 40.14 51.25 57.03 54.69
T5 @ 16k 21.4 22.0 21.24 57.62 55.18
DEPTH @ 16k 36.47 40.33 53.13 57.52 55.44
T5 @ 32k 21.75 21.75 20.63 57.62 56.1
DEPTH @ 32k 38.04 44.26 54.05 58.5 57.37
T5 @ 64k 21.92 21.53 21.24 57.23 57.01
DEPTH @ 64k 42.24 45.85 55.96 60.94 72.58
T5 @ 128k 21.09 33.96 42.63 60.16 60.01
DEPTH @ 128k 45.0 47.71 59.57 64.65 78.0
T5 @ 256k 22.85 37.92 44.85 58.89 61.91
DEPTH @ 256k 48.68 48.93 61.33 65.33 81.69
T5 @ 512k 26.61 41.97 52.29 61.33 60.89
DEPTH @ 512k 52.59 51.66 65.92 66.85 83.81
T5 @ 1M 28.54 43.22 50.98 61.13 65.48
DEPTH @ 1M 52.39 50.07 63.89 67.92 84.52
Table 6: DiscoEval Downstream Continuous Pre-Training (CPT) Results
Model SP DC
Arxiv Wiki Rocstory Wiki Chat
T5 @ 2k 62.26 51.9 74.83 92.09 71.44
DEPTH @ 2k 44.14 49.34 74.78 89.99 70.46
T5 @ 4k 61.33 52.39 74.98 91.72 74.27
DEPTH @ 4k 56.62 51.2 67.65 91.46 72.02
T5 @ 8k 63.63 52.03 77.0 92.33 73.1
DEPTH @ 8k 55.03 51.56 71.02 91.99 71.97
T5 @ 16k 60.94 51.78 76.49 91.53 74.71
DEPTH @ 16k 58.86 52.08 76.15 92.63 72.71
T5 @ 32k 60.69 52.22 75.61 92.33 74.02
DEPTH @ 32k 59.35 53.27 78.08 92.48 72.85
T5 @ 64k 58.96 52.25 76.07 92.58 73.73
DEPTH @ 64k 58.57 52.95 76.1 92.58 73.24
T5 @ 128k 60.13 51.03 75.76 91.14 73.44
DEPTH @ 128k 70.56 54.77 82.42 92.53 73.96
T5 @ 256k 59.03 52.66 66.77 92.31 72.22
DEPTH @ 256k 67.07 53.0 76.71 92.07 72.9
Appendix G NI Results
In the from-scratch setting (Figure 6(a) and table 7), we observe that DEPTH outperforms T5 in the NI benchmark, with a notable leap in performance between steps 16k and 32k. This indicates that DEPTH’s pre-training objective is more effective at learning representations that are beneficial for the NI task. However, at steps 2k, 8k, and 16k, DEPTH underperforms compared to T5, suggesting that the benefits of DEPTH’s pre-training objective may not be immediately apparent in the early stages of training.
However, in the continuously pre-trained setting (Figure 6(b) and table 8), we find that DEPTH’s pre-training harms downstream performance compared to T5. Additionally, we observe that CPT models are less sensitive to learning rate and can train effectively across a wider range of learning rates, with the exception of DEPTH in the early stages of CPT, where it is adapting to a task that differs from its initial pre-training. This robustness to learning rate is a positive property that the FS models did not exhibit, likely due to limitations in training scale (e.g., small batch size, avoiding packing, and training on fewer tokens overall). Furthermore, early in the CPT process, DEPTH’s performance is somewhat unstable, possibly due to the domain shift from T5’s pre-training task to DEPTH’s pre-training task. Interestingly, lower learning rates perform worse for DEPTH after CPT, suggesting that the model needs to adjust its representations more substantially to adapt to the downstream task.
Taken as a whole, while DEPTH performs well on NLU tasks (e.g, DC and SP from DiscoEval), when applied with limited pre-training and no supervised fine-tuning [60], its performance may be weaker than other encoder-decoders with respect to downstream tasks involving long-form text generation.
Table 7: NI benchmark results for From Scratch (FS) pre-training. The first row reports the performance of the baseline T5 model, while all later rows show the best reported results on the validation set across 3 attempted learning rates.
Model Step RougeL
Baseline T5 1M 42.48
T5 2k 8.36
DEPTH 2k 10.92
T5 4k 10.11
DEPTH 4k 11.03
T5 8k 10.15
DEPTH 8k 9.06
T5 16k 10.82
DEPTH 16k 10.43
T5 32k 10.68
DEPTH 32k 23.51
T5 64k 12.89
DEPTH 64k 30.23
T5 128k 18.24
DEPTH 128k 32.24
T5 256k 26.36
DEPTH 256k 32.63
T5 512k 28.05
DEPTH 512k 34.72
T5 1M 29.6
DEPTH 1M 33.8
Table 8: NI benchmark results for Continuous Pre-Training (CPT). All rows show the best reported results on the validation set across 3 attempted learning rates.
Model Step NI RougeL
T5 2k 41.96
DEPTH 2k 39.15
T5 4k 42.72
DEPTH 4k 37.85
T5 8k 42.83
DEPTH 8k 38.04
T5 16k 42.88
DEPTH 16k 38.19
T5 32k 43.56
DEPTH 32k 37.79
T5 64k 43.06
DEPTH 64k 38.99
T5 128k 42.58
DEPTH 128k 39.19
T5 256k 43.29
DEPTH 256k 37.86